(a) On January 1, 2020, Metlock Inc. purchased land that had an assessed value of $316,000 at the time of purchase. A $518,000, zero-interest-bearing note due January 1, 2023, was given in exchange. There was no established exchange price for the land, nor a ready fair value for the note. The interest rate charged on a note of this type is 12%.

Determine at what amount the land should be recorded at January 1, 2020, and the interest expense to be reported in 2020 related to this transaction. (Round answers to 0 decimal places, e.g. 38,548.)

Land to be recorded at January 1, 2020 $
Interest expense to be reported $


(b) On January 1, 2020, Bonita Furniture borrowed $6,100,000 (face value) from Sinise Co., a major customer, through a zero-interest-bearing note due in 4 years. Because the note was zero-interest-bearing, Bonita Furniture agreed to sell furniture to this customer at lower than market price. A 10% rate of interest is normally charged on this type of loan.

Prepare the journal entry to record this transaction. (Round answers to 0 decimal places, e.g. 38,548. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually.)

Account Titles and Explanation

Debit

Credit


Determine the amount of interest expense to report for 2020. (Round answer to 0 decimal places, e.g. 38,548.)

Interest expense to be reported for 2020 $
Answer #1

(a)

Land price is equal to present value of note given in exchange.

Land value to be recorded on January 1, 2020 = $518,000 / (1+r)^3 = $518,000 / (1+0.12)^3 = $368,702.

Interest expense for 2020 = Land value X interest rate = $368,702 X 12% = $44,244

(b)

Note carrying amount = $6,100,000/(1+10%)^4 = $4,166,382

Interest for the year = $4,166,382 X 10% = $416,638

Below are journal entries:

Account title Debit Credit
Cash $     4,166,382
Note payable $     4,166,382
[Note written in exchange for cash]
Interest expense $ 416,638.20
Note payable $        416,638
[Interest accrued for the year]